# Gradients: When Markets Meet Fine-tuning - A Distributed Approach to Model Optimisation

### v1

@@ 3.5 Miner Scoring
$S_{\text{task}}(m_{i})=\begin{cases}S_{\text{first}}&\text{if rank}(m_{i})=1\\ S_{\text{penalty}}&\text{if rank}(m_{i})>|\mathcal{M}_{\text{valid}}|\cdot(1-\rho_{\text{penalty}})\\ 0&\text{otherwise}\end{cases}$ (11)
@@
$S_{\text{adjusted}}(m_{i},\mathcal{T})=S_{\text{task}}(m_{i})\cdot W_{\text{task}}(\mathcal{T})$ (12)
@@
$S_{\text{temporal}}(m_{i})=\sum_{w\in W}\omega_{w}\cdot\text{Normalize}\left(\sum_{\mathcal{T}\in\mathcal{T}w}S{\text{adjusted}}(m_{i},\mathcal{T})\right)$ (13)
where $W$ is the set of time windows 1-day, 3-day, 7-day, $\omega_{w}$ is the weight for window $w$ (0.3, 0.3, and 0.4 respectively), and $\mathcal{T}_{w}$ is the set of tasks in window $w$. The final score undergoes a non-linear transformation to create appropriate reward differentials:
$S_{\text{final}}(m_{i})=\beta\cdot S_{\text{sigmoid}}(m_{i})+\omega\cdot x_{i}$ (14)
@@
$S_{\text{sigmoid}}(m_{i})$ $=\left(\frac{1}{1+\exp(-\gamma\cdot(x_{i}-\mu))}\right)^{\nu}$ (15)
and $x_{i}=\frac{\text{quality\_score}_{i}}{\text{max\_score}}$ is the normalised input score, with parameters $\beta=0.7$ (sigmoid weight), $\omega=0.05$ (linear weight), $\gamma=9$ (sigmoid steepness), $\mu=0.5$ (sigmoid shift), and $\nu=0.75$ (sigmoid power). This transformation creates pronounced rewards for top performers while maintaining reasonable incentives for all participants.
3.6 Blockchain Integration and Economic Incentives
@@
3.7 System Fault Tolerance
@@
4 Methodology
4.1 Experimental Design
We conducted 180 controlled experiments to evaluate the effectiveness of decentralised hyperparameter optimisation across diverse models, datasets, and task types. To ensure fair comparison, all experiments used identical conditions across platforms: the same base models, datasets, training time allocations, and evaluation metrics. We trained all models using the Gradients platform’s distributed fine-tuning approach, then evaluated them using task-specific metrics against baseline implementations from leading AutoML services using these identical experimental conditions. We’re focussing the experimental space on instruct and diffusion tasks - since there are no other platforms to reliably compare DPO and GRPO tasks.
4.2 Model Selection
We tested five model families spanning three orders of magnitude in size. EleutherAI Pythia-70M [1] represents the smallest scale, testing extreme parameter efficiency. The Qwen2 family [39] (1.5B, 7B) provides popular open models with strong multilingual capabilities. Llama-3.1-8B and Llama-3.3-70B [16, 25] are Meta’s latest instruction-tuned models. DeepSeek-R1-Distill-70B [6] represents state-of-the-art reasoning-optimised architecture.
For diffusion models, we evaluated two state-of-the-art text-to-image architectures: SDXL 1.0 [30] (Stable Diffusion XL), the workhorse of open image generation, and Flux.1 [22], a next-generation architecture with improved coherence.
4.3 Training Protocol
@@
4.4 Evaluation Methodology
4.4.1 Language Model Evaluation
@@
4.4.2 Diffusion Model Evaluation
@@
$L_{\text{text}}=\frac{1}{N}\sum_{i=1}^{N}\|I_{\text{original},i}-G(I_{\text{noisy},i},p_{i})\|_{2}^{2}$ (16)
@@
$L_{\text{no-text}}=\frac{1}{N}\sum_{i=1}^{N}\|I_{\text{original},i}-G(I_{\text{noisy},i},\emptyset)\|_{2}^{2}$ (17)
@@ 4.5 Baseline Comparisons
HuggingFace AutoTrain [37]: State-of-the-art AutoML with Bayesian hyperparameter optimisation
@@
For diffusion models, we compared against CivitAI [38], the only competitor available for this form of lora training.
4.6 Language Model Tasks
@@
Mathematical Reasoning: GSM8K [4] (grade school maths), OpenThoughts-114k [36] (step-by-step solutions), and verified mathematical solutions with chain-of-thought reasoning.
@@
Code Generation: CodeAlpaca-20k [3] (instruction-following), self-instruct-starcoder [5] (compilation tasks), and CodeFeedback-Filtered (complex programming challenges).
Retrieval-Augmented Generation: RAG-Instruct [14] (document-grounded QA), Ruler [18] (long-context retrieval), and RAGBench [35, 15] (multi-hop reasoning).
@@
4.7 Diffusion Model Tasks
@@ 5 Results
5.1 Task-Specific Performance Analysis
The effectiveness of decentralised optimisation varies by task complexity. Figure 1 shows comprehensive statistical breakdown across task types. RAG and reasoning tasks—requiring complex contextual understanding—show the largest gains, while well-studied translation tasks show more modest but consistent improvements.
5.2 Overall Performance and Relative Analysis
@@
Figure 2: Win rates against leading AutoML platforms. Matrix shows percentage of experiments where Gradients achieved lower loss than competitors across 180 controlled experiments.
@@
Figure 3: Performance magnitude analysis. (a) Improvement distributions across all experiments demonstrate both consistency and magnitude of gains over competitors. (b) Spider chart visualises performance relative to cross-provider median, where positive values indicate above-median performance.
5.3 Model Scale Effects
A surprising finding emerges from model scale analysis (Figure 4): mid-scale models (7-8B parameters) benefit most from hyperparameter optimisation. This challenges conventional wisdom about model robustness and suggests an optimal complexity range for hyperparameter sensitivity, although we wish to note to the reader that there were fewer 70b model experiments due to prohibitive costs (over \$10,000 per experiment for GCP 70b llama for example) - so experiment noise could play a role.
@@ 5.3 Model Scale Effects
Figure 4: Scale and architecture effects. (a) Performance improvements peak at 7-8B parameters, suggesting optimal hyperparameter sensitivity. Very small (70M) and large (70B) models show reduced gains. (b) Diffusion model improvements increase with task difficulty - Gradients’ advantage grows substantially as noise levels rise, demonstrating superior optimization for challenging denoising scenarios.
5.4 Implications and Insights
Our results reveal two potential insights about hyperparameter optimisation:
1. The optimisation landscape is more complex than assumed. The 100% win rate against TogetherAI and 82.8% against HuggingFace—across diverse tasks and models—indicates that standard configurations occupy local optima. The broad improvement distributions suggest multiple superior configurations exist but remain undiscovered by conventional search.
2. Task complexity determines optimisation potential. Complex tasks (RAG, reasoning) show 30-40% improvements while mature tasks (translation) show 10-20%. This pattern suggests that as we tackle increasingly sophisticated AI challenges, the value of the decentralised mining pool seemingly grows proportionally.
These findings validate our core thesis: decentralised competitive search, by aligning economic incentives with exploration, systematically discovers configurations that centralised approaches miss. The consistency across 180 experiments provides strong evidence that current fine-tuning practices leave substantial performance unrealised.
6 Discussion
Why Competitive Approaches Work Our results suggest that competition alters how hyperparameter optimisation occurs. In our experiments, centralised systems followed predetermined search strategies, whilst independent miners developed varied approaches. The tournament structure in our system provided incentives for exploring configurations that centralised methods did not test, as evidenced by the consistent performance improvements across task types. The 82.8% win rate against HuggingFace and 100% against other providers indicates that competitive exploration can identify configurations in regions of hyperparameter space that standard AutoML approaches do not examine.
Task Complexity Correlation We observed a pattern where certain complex tasks showed larger performance improvements: RAG tasks improved by 23-48 and reasoning tasks by 24-31, whilst translation tasks improved by 9-57 and mathematical tasks by 1-26. This suggests that the type of complexity matters—tasks requiring multi-step reasoning and information synthesis showed the most substantial gains, while mathematical problem-solving showed more modest improvements despite its inherent complexity.
Computational Requirements One limitation with the Gradients approach is that it requires substantially more computational resources than centralised alternatives. Each task involves 8-15 parallel training runs instead of a single optimisation attempt, resulting in proportionally higher energy consumption and cost. Whilst this produces better-performing models, the environmental and economic costs may limit applicability in resource-constrained settings. The energy trade-off between optimisation quality and computational efficiency represents a fundamental limitation of the competitive approach.
7 Conclusion
@@
Our findings reveal that hyperparameter optimisation contains unexploited potential that centralised search strategies often miss. Through competitive economic coordination, Gradients’ discovers superior fine-tuning configurations across diverse foundation model architectures and task types, achieving 82.8% win rate against HuggingFace AutoTrain and perfect performance against other leading platforms. The 30-40% improvements on complex reasoning tasks and 23% gains for diffusion models indicate that current AutoML approaches explore only a fraction of viable hyperparameter space.

### v2

@@ 3.5. Miner Scoring
(9) $S_{\text{task}}(m_{i})=\begin{cases}S_{\text{first}}&\text{if rank}(m_{i})=1\\ S_{\text{penalty}}&\text{if rank}(m_{i})>|\mathcal{M}_{\text{valid}}|\cdot(1-\rho_{\text{penalty}})\\ 0&\text{otherwise}\end{cases}$
@@
(10) $S_{\text{adjusted}}(m_{i},\mathcal{T})=S_{\text{task}}(m_{i})\cdot W_{\text{task}}(\mathcal{T})$
@@
(11) $S_{\text{temporal}}(m_{i})=\sum_{w\in W}\omega_{w}\cdot\text{Normalize}\left(\sum_{\mathcal{T}\in\mathcal{T}w}S{\text{adjusted}}(m_{i},\mathcal{T})\right)$
where $W$ is the set of time windows {1-day, 3-day, 7-day}, $\omega_{w}$ is the weight for window $w$ (0.3, 0.3, and 0.4 respectively), and $\mathcal{T}_{w}$ is the set of tasks in window $w$.
The final score undergoes a non-linear transformation to create appropriate reward differentials:
(12) $S_{\text{final}}(m_{i})=\beta\cdot S_{\text{sigmoid}}(m_{i})+\omega\cdot x_{i}$
@@
(13) $S_{\text{sigmoid}}(m_{i})$ $=\left(\frac{1}{1+\exp(-\gamma\cdot(x_{i}-\mu))}\right)^{\nu}$
and $x_{i}=\frac{\text{quality\_score}_{i}}{\text{max\_score}}$ is the normalised input score, with parameters $\beta=0.7$ (sigmoid weight), $\omega=0.05$ (linear weight), $\gamma=9$ (sigmoid steepness), $\mu=0.5$ (sigmoid shift), and $\nu=0.75$ (sigmoid power).
This transformation creates pronounced rewards for top performers while maintaining reasonable incentives for all participants.
3.6. Blockchain Integration and Economic Incentives
@@
3.7. System Fault Tolerance
@@
4. Methodology
4.1. Experimental Design
We conducted 180 controlled experiments to evaluate the effectiveness of decentralised hyperparameter optimisation across diverse models, datasets, and task types. To ensure fair comparison, all experiments used identical conditions across platforms: the same base models, datasets, training time allocations, and evaluation metrics. We trained all models using the Gradients platform’s distributed fine-tuning approach, then evaluated them using task-specific metrics against baseline implementations from leading AutoML services using these identical experimental conditions. We focus the experimental space on instruct and diffusion tasks - since there are no other platforms to reliably compare DPO and GRPO tasks.
4.2. Model Selection
We tested five model families spanning three orders of magnitude in size. EleutherAI Pythia-70M (Biderman et al., 2023) represents the smallest scale, testing extreme parameter efficiency. The Qwen2 family (Yang et al., 2024) (1.5B, 7B) provides popular open models with strong multilingual capabilities. Llama-3.1-8B and Llama-3.3-70B (Grattafiori et al., 2024; Meta, 2024) are Meta’s latest instruction-tuned models. DeepSeek-R1-Distill-70B (DeepSeek-AI et al., 2025) represents state-of-the-art reasoning-optimised architecture.
For diffusion models, we evaluated two state-of-the-art text-to-image architectures: SDXL 1.0 (Podell et al., 2023) (Stable Diffusion XL), the workhorse of open image generation, and Flux.1 (Labs, 2024), a next-generation architecture with improved coherence.
4.3. Training Protocol
@@
4.4. Evaluation Methodology
4.4.1. Language Model Evaluation
@@
4.4.2. Diffusion Model Evaluation
@@
(14) $L_{\text{text}}=\frac{1}{N}\sum_{i=1}^{N}\|I_{\text{original},i}-G(I_{\text{noisy},i},p_{i})\|_{2}^{2}$
@@
(15) $L_{\text{no-text}}=\frac{1}{N}\sum_{i=1}^{N}\|I_{\text{original},i}-G(I_{\text{noisy},i},\emptyset)\|_{2}^{2}$
@@ 4.5. Baseline Comparisons
HuggingFace AutoTrain (Thakur, 2024): State-of-the-art AutoML with Bayesian hyperparameter optimisation
@@
For diffusion models, we compared against CivitAI (Wei et al., 2024), the only competitor available for this form of lora training.
4.6. Language Model Tasks
@@
Mathematical Reasoning: GSM8K (Cobbe et al., 2021) (grade school maths), OpenThoughts-114k (Team, 2025) (step-by-step solutions), and verified mathematical solutions with chain-of-thought reasoning.
@@
Code Generation: CodeAlpaca-20k (Chaudhary, 2023) (instruction-following), self-instruct-starcoder (CodeParrot, 2023) (compilation tasks), and CodeFeedback-Filtered (complex programming challenges).
Retrieval-Augmented Generation: RAG-Instruct (FreedomIntelligence, 2024) (document-grounded QA), Ruler (Hsieh et al., 2024) (long-context retrieval), and RAGBench (Tang and Yang, 2024; Friel et al., 2024) (multi-hop reasoning).
@@
4.7. Diffusion Model Tasks
@@ 5. Results
The effectiveness of decentralised optimisation varies by task complexity. Figure 1 shows breakdown across task types. RAG and reasoning tasks—requiring complex contextual understanding—show the largest gains, while math show more modest, but consistent improvements.
@@
Figure 2. Win rates against leading AutoML platforms. Matrix shows percentage of experiments where Gradients achieved lower loss than competitors across 180 controlled experiments.
@@
Figure 3. Performance magnitude analysis. (a) Improvement distributions across all experiments demonstrate both consistency and magnitude of gains over competitors. (b) Spider chart visualises performance relative to cross-provider median, where positive values indicate above-median performance.
5.1. Model Scale Effects
Figure 4 shows the relationship between model scale and performance improvements. We observe that mid-scale models (7-8B parameters) achieve the largest gains from competitive hyperparameter optimisation, while both smaller (70M) and larger (70B) models show more modest improvements. For diffusion models, we evaluated performance across varying noise levels (0.1 to 0.9 denoise strength). The competitive advantage of Gradients increases as the denoising task becomes more challenging—the performance gap widens substantially at higher noise levels, demonstrating that the distributed approach excels particularly when the reconstruction task is most difficult. Note that the number of experiments for 70B models was limited due to computational costs (exceeding \$10,000 per training run on some platforms), which may affect the statistical significance of results at that scale.
@@ 5.1. Model Scale Effects
Figure 4. Scale and architecture effects. (a) Performance improvements peak at 7-8B parameters, suggesting optimal hyperparameter sensitivity. Very small (70M) and large (70B) models show reduced gains. (b) Diffusion model improvements increase with task difficulty - Gradients’ advantage grows substantially as noise levels rise, demonstrating superior optimization for challenging denoising scenarios.
6. Discussion
Why Competition Discovers Better Configurations: Our results suggest three mechanisms drive competitive advantage in hyperparameter optimization. First, centralized search explores limited regions of configuration space—the consistency of our wins indicates that standard AutoML confines itself to local optima while miners discover superior global configurations. Second, the tournament structure fundamentally alters exploration incentives: miners must outperform competitors rather than achieve merely acceptable performance, naturally driving search into unexplored regions. Third, strategy opacity prevents convergence—because miners keep configurations proprietary, the system maintains diversity rather than collapsing to common approaches.
Adaptation Without Central Coordination: Traditional AutoML platforms improve through explicit algorithmic updates—a slow cycle from research to deployment. Gradients evolves continuously through market dynamics. During our experimental period, we observed progressive sophistication: early submissions used standard techniques while later submissions incorporated adaptive schedules and novel lora techniques. This distributed innovation emerges from economic pressure alone.
Computational Trade-offs:
The primary limitation of competitive optimization is resource consumption. Each task requires 8-25 parallel training runs versus single runs in centralized systems, proportionally increasing energy use and cost. Whether 20-40% performance gains justify this overhead depends on application requirements. High-stakes deployments may warrant the expense, while routine fine-tuning may not. Future work should explore adaptive competition—varying miner pools based on task novelty and importance—to balance discovery with efficiency.
Strategic Opacity and Emergent Optimization:
Unlike traditional AutoML where optimization paths are internally observable, Gradients operates as a black box we observe superior performance without visibility into specific winning configurations. This opacity is not a design flaw but an essential feature that drives discovery. Strategy secrecy preserves competitive advantage, incentivizing miners to explore novel combinations rather than converge on known parameters. This mirrors biological evolution and financial markets, where complex adaptive behavior emerges without central coordination.
Despite configuration opacity, output artifacts reveal strategic patterns that contradict theoretical guidance. Miners consistently employ rsLoRA with ranks of 64-128, despite theory suggesting rank 1 should suffice (Hu et al., 2022). They select rsLoRA for small datasets but prefer full fine-tuning for larger corpuses—empirical discoveries that conventional AutoML might never explore due to theoretical priors. This creates a paradox: the system’s effectiveness may depend on its inscrutability. Traditional AutoML achieves reproducibility through transparency but limits search to theoretically justified strategies. Gradients discovers superior configurations precisely because miners explore regions that gradient-based or Bayesian optimization would never sample—using ranks orders of magnitude higher than theory suggests or learning rates far outside conventional ranges.
Accepting emergent complexity may be necessary for optimal hyperparameter optimization. Future work could explore controlled transparency mechanisms—periodic strategy disclosure after competitive advantage windows or differential privacy for configuration sharing—but maintaining some degree of strategic opacity appears crucial for driving continuous innovation.
7. Conclusion
@@
Our findings reveal that hyperparameter optimisation contains unexploited potential that centralised search strategies often miss. Through competitive economic coordination, Gradients discovers superior fine-tuning configurations across diverse foundation model architectures and task types, achieving 82.8% win rate against HuggingFace AutoTrain and perfect performance against other leading platforms. The 30-40% improvements on complex reasoning tasks and 23% gains for diffusion models indicate that current AutoML approaches explore only a fraction of viable hyperparameter space.